OpenStudy (anonymous):

prove that sin(n+1)x sin(n+2)x + cos(n+1)x cos(n+2)x =cosx

OpenStudy (anonymous):

let (n+1)x be A and (n+2)x be B then apply the formula cos(a-b) = sin a sin b + cos a cos b by doing that you'll get cos[nx +x-nx-2x] = cos x since cos(-x) = cos x therefore cos x = cos x

OpenStudy (anonymous):

I'm going to break this down into two parts and then add them: The first part: \[\sin[(n+1)x]\sin[(n+2)x]\] use the identity: \[\sin \alpha \sin \beta=\frac{1}{2}[\cos(\alpha-b)-\cos (\alpha+\beta)]\] in our case: \[\alpha= (n=1)x, \beta=(n+2)x\] Thus: \[\sin[(n+1)x]\sin[(n+2)x]=\] \[\frac{1}{2}[\cos[(n+1)x-(n+2)x]-\cos[(n+1)x+(n+2)x]]\] \[=\frac{1}{2}[\cos(-x)-\cos(2nx+3x)]\] The second part: \[\cos[(n+1)x]\cos[(n+2)x]\] use the identity: \[\cos \alpha \cos \beta=\frac{1}{2}[\cos(\alpha-\beta)+\cos(\alpha+\beta)]\] in our case: \[\alpha=(n+1)x, \beta=(n+2)x\] Thus: \[\cos[(n+1)x]\cos[(n+2)x]=\] \[\frac{1}{2}[\cos[(n+1)x-(n+2)x]+\cos[(n+1)x+(n+2)x]]\] \[=\frac{1}{2}[\cos(-x)+\cos(2nx+3x)]\] Now add parts one and two: \[\frac{1}{2}[\cos(-x)-\cos(2nx+3x)]+\frac{1}{2}[\cos(-x)+\cos(2nx+3x)]=\] \[\frac{1}{2}cos(-x)-\frac{1}{2}\cos(2nx+3x)+\frac{1}{2}cos(-x)+\frac{1}{2}\cos(2nx+3x)=\] \[\frac{1}{2}\cos(-x)+\frac{1}{2}\cos(-x)=\cos(-x)\] Since cosine is an even function: cos(-x)=cos(x)